OpenStudy (anonymous):

Write an equation of the circle with center (5,−2) and radius 4

OpenStudy (mrnood):

so you must be studying the equations of circles the general form is (x-a)^2 +(y-b)^2 = r^2 where (a,b) is the coordinate of the centre, and r is the radius in oyur case a=5 b=-2 r=4 so pput those values into the general form to get your answer. (hint - be careful of the - signs...)

OpenStudy (jdoe0001):

hint: \(\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ 5}},{\color{blue}{ -2}})\qquad radius={\color{purple}{ 4}}\)

OpenStudy (anonymous):

B is negative -2 how would i put that?

OpenStudy (mrnood):

it is still (y-b) so put the -2 in place of b (that leaves 2 - signs in front of the 2)

OpenStudy (anonymous):

x-5^2+y--2^2=4^2???

OpenStudy (mrnood):

yes - but you need the brackets to make it clear you can simplify the two minuses you can work out r^2